Find the value of x:
3x-2 = 10-x
Use either the multiplicative inverse or the additive inverse.

Answer by ewatrrr(24785)

 
Hi,
Find the value of x (Isolating x on one side of the Equation)
3x-2 = 10-x || Using additive inverse of -x {Adding x to both sides)
4x - 2 = 10 || Using additive inverse of -2 (Adding 2 to both sides)
4x = 12 || Using multiplicative inverse of 4 (multiplying both sides by 1/4)
Yes! x = 3